# UHECR deflections in the Galactic magnetic field

### v1

@@ I Introduction
Two modern ultra-high-energy cosmic ray (UHECR) experiments, the Pierre Auger Observatory (Auger) [5] and the Telescope Array (TA) [6, 7], were largely motivated by the possibility of UHECR astronomy. At that time, it was believed that most of the UHECR above the ankle at $E>3$ EeV were protons [8]. In this case, the deflections of UHECR with $E>60$ EeV in the Galactic magnetic field (GMF) with the magnitude of $B_{\mathrm{Gal}}\sim\mu$G were expected to be on the order of a few degrees [9]. The extragalactic field was not expected to produce large deflections either. Indeed, since protons lose energy in interactions with the cosmic microwave background (CMB) radiation [10, 11], their propagation distance is limited to $50-100$ Mpc. Taking into account constraints on the intergalactic magnetic field (IGMF) $B_{\mathrm{IGMF}}<1$ nG, the latter would not significantly change the trajectories of UHECR protons on their way from sources to our Galaxy. As a result, most UHECR protons with $E>60$ EeV should point back to their sources within a few degrees, and one could try to identify the sources after collecting a few years of data with the Auger and TA experiments. Moreover, exact modeling of the GMF would not be required if one could identify the sets of events of different energies coming from the same source (see, e.g., [12, 13]).
@@ II GMF model dependence of UHECR deflections
The GMF is an unavoidable medium that deflects charged UHECRs, thereby obscuring their sources. While the general structure of the GMF is consistent across recent models, the parametrization and exact profiles of its components differ, making the UHECR deflection highly model-dependent. In this section, we compare the predicted UHECR deflections in popular GMF models, including the two most recent models, UF23 and KST24. We discuss the GMF components responsible for UHECR deflections in certain regions of the sky, as well as the modeling uncertainties associated with these components and possible future improvements. Our analysis is limited to the coherent GMF only; the turbulent GMF is not considered.
@@ II.1 Deflection magnitudes
Figure 1: Angular deflections of $R=20$ EV (upper panels) and $R=10$ EV (lower panels) particles in PT11, JF12, UF23 base and KST24 GMF models. The position of the dipole is for the energies $E^{\mathrm{TA}}_{\mathrm{Auger}}\geq{}^{10~{}\mathrm{Eev}}_{8.55~{}\mathrm{Eev}}$ from [20]. The skymaps are shown in the Galactic coordinates using the Mollweide projection. The Galactic center is at the center of the maps and longitude increases to the left.
@@
The total difference shown in Figs. 1 and 2 accumulates effects of various origins. The first effect can be attributed to the different nature of the magnetic field tracers used to construct the GMF model. The most pronounced effect of this type is visible in the central part of the Galaxy ($-60^{\circ}<l<60^{\circ}$). The PT11 model predicts smaller deflections in this region of the sky compared to all other models. This is because PT11 was fitted only to the rotation measures (RMs) of extragalactic sources, whereas three other models were also fitted to synchrotron data from the WMAP/Planck. RM data is sensitive to the magnetic field parallel to the line of sight (LOS) and is thus almost insensitive to the X-field component of the GMF, which crosses the Galactic plane from below upwards in the central part of the Galaxy and is mainly perpendicular to the LOS. The presence of the X-field component is clearly identifiable in the synchrotron data and is primarily responsible for the large deflections observed in the central Galaxy at all latitudes in JF12, UF23 and KST24.
@@
The last type of uncertainty is related to the interpretation of the data itself. Namely, the local features appearing in the data should be masked or modeled independently, while the large-scale features, associated with the coherent GMF, should be included in the model. The main obstacle here is that information about the distance to magnetic structures is still very limited and available only for a few of them. The interpretation of the data is one of the primary sources (along with parametrizations) of the differences between the most recent models, UF23 and KST24. While both were fitted to very similar datasets  [^2] UF23 and KST24 were fitted to the latest compilation of the RMs of extragalactic sources [2023ApJS..267...28V] and polarized synchrotron data from WMAP/Planck at 23/30 GHz [WMAP_23GHz], the predicted UHECR deflections differ quite substantially in a significant fraction of the sky.
To better understand the differences between UF23 and KST24 models, it is useful to divide the sky into three regions: the central Galaxy ($-60^{\circ}<l<60^{\circ}$), the outer Galaxy at high latitudes, and the outer Galaxy at low latitudes. In the central Galaxy the deflections are controlled by the X-field and its interplay with the toroidal field of the halo (the GMF halo component parallel to the disk). As shown in the Fig. 1 the deflections in this region form a ”butterfly-like” pattern. The typical deflections are large, exceeding $>20^{\circ}$ for $20$ EV particles even far above and below the Galactic plane. This is a robust prediction of the recent models JF12, UF23, and KST24 all of which take into account the polarized synchrotron data.
@@
In contrast, in the outer Galaxy at high latitudes ($|b|\gtrsim 20^{\circ}$), all models predict relatively small deflections, likely not exceeding $30^{\circ}$. Currently, there is no data suggesting the presence of a strong GMF in these directions. Interestingly, the TA hotspot, marked by the black circle in Fig. 1, is located in this region. As a result, UHECR source searches in this area are likely to be less dependent on the specific GMF model.
@@ II.2 Deflection directions
In Fig. 5, we present the Amaterasu source localization. The purple contour for the KST24 model was obtained by varying the model parameters, as well as particle energy, within their one-sigma uncertainty. Additionally, $5^{\circ}$ smearing by a toy model of the turbulent GMF was taken into account, following the prescription from [45]. For the KST24 contour, the Amaterasu particle was assumed to be an iron nucleus. The contour size is primarily driven by the observed energy uncertainty. This, however, does not include “systematic” uncertainties related to modeling of GMF, so the actual uncertainties are larger. This is illustrated by the green contour in Fig. 5, which shows the source localization from a Monte Carlo analysis of [42] for the UF23 models. Note that, for both the green contour from [42] and the purple contour presented in this paper for KST24, the particles’ energy was rescaled to $212$ EeV to account for systematic bias in energy reconstruction for heavy primaries. Note also that the observed direction of the Amaterasu event points to the Loop I, the bright feature in the polarized synchrotron data (shown with black dashed lines in Fig. 5). Recently, it was shown to be a Galactic-scale outflow [46, 47], but it is not included in any GMF model. Incorporating this feature into GMF models could influence the reconstructed direction.
@@ III Discussion and Summary
In summary, in most directions on the sky, UHECR deflections are still too uncertain — both in direction and magnitude — to enable source identification through backtracking. A notable exception is the region at high latitudes, roughly at $30^{\circ}<b<70^{\circ}$ in the outer Galaxy, where all existing models consistently predict relatively small deflections. Interestingly, this is the same region where the TA hotspot has been observed. Regarding the Amaterasu particle, the precise localization of its source cannot be achieved with the KST24 model, although the reconstructed source region is close to that found with the UF23 model.

### v2

# UHECR deflections in the coherent Galactic magnetic field

@@ I Introduction
[…]
@@ II GMF model dependence of UHECR deflections
The GMF is an unavoidable medium that deflects charged UHECRs, thereby obscuring their sources. While the general structure of the GMF is consistent across recent models, the parametrization and exact profiles of its components differ, making the UHECR deflection highly model-dependent. In this section, we compare the predicted UHECR deflections in popular coherent GMF models, including the two most recent models, UF23 and KST24. We discuss the GMF components responsible for UHECR deflections in certain regions of the sky, as well as the modeling uncertainties associated with these components and possible future improvements. Our analysis is focused on the coherent GMF; however, deflections by the turbulent GMF are briefly discussed in Section II.3.
@@ II.1 Deflection magnitudes
Figure 1: Angular deflections of $R=20$ EV (upper panels) and $R=10$ EV (lower panels) particles in PT11, JF12, UF23 base and KST24 models of the coherent GMF. The position of the dipole is for the energies $E^{\mathrm{TA}}_{\mathrm{Auger}}\geq{}^{10~\mathrm{Eev}}_{8.55~\mathrm{Eev}}$ from [20]. The skymaps are shown in the Galactic coordinates using the Mollweide projection. The Galactic center is at the center of the maps and longitude increases to the left. Note that all small-scale features visible on the maps will be washed out by the turbulent GMF, see Sec. II.3 for details.
@@
The total difference shown in Figs. 1 and 2 accumulates effects of various origins. The first effect can be attributed to the different nature of the magnetic field tracers used to construct the GMF model. The most pronounced effect of this type is visible in the central part of the Galaxy ($60^{\circ}<l<60^{\circ}$). The PT11 model predicts smaller deflections in this region of the sky compared to all other models. This is because PT11 was fitted only to the rotation measures (RMs) of extragalactic sources, whereas three other models were also fitted to synchrotron data from the WMAP/Planck. RM data is sensitive to the magnetic field parallel to the line of sight (LOS) and is thus almost insensitive to the X-field component of the GMF, which crosses the Galactic plane from below upwards in the central part of the Galaxy and is mainly perpendicular to the LOS. The presence of the X-field component is clearly identifiable in the synchrotron data and is primarily responsible for the large deflections observed in the central Galaxy at all latitudes in JF12, UF23 and KST24.
@@
The last type of uncertainty is related to the interpretation of the data itself. Namely, the local features appearing in the data should be masked or modeled independently, while the large-scale features, associated with the coherent GMF, should be included in the model. The main obstacle here is that information about the distance to magnetic structures is still very limited and available only for a few of them. The interpretation of the data is one of the primary sources (along with parametrizations) of the differences between the most recent models, UF23 and KST24. While both were fitted to very similar datasets (UF23 and KST24 were fitted to the latest compilation of the RMs of extragalactic sources [37] and polarized synchrotron data from WMAP/Planck at 23/30 GHz [38]), the predicted UHECR deflections differ quite substantially in a significant fraction of the sky.
To better understand the differences between UF23 and KST24 models, it is useful to divide the sky into three regions: the central Galaxy ($60^{\circ}<l<60^{\circ}$), the outer Galaxy at high latitudes, and the outer Galaxy at low latitudes. In the central Galaxy the deflections are controlled by the X-field and its interplay with the toroidal field of the halo (the GMF halo component parallel to the disk). As shown in the Fig. 1 the deflections in this region form a ”butterfly-like” pattern. The typical deflections are large, exceeding $20^{\circ}$ for $20$ EV particles even far above and below the Galactic plane. This is a robust prediction of the recent models JF12, UF23, and KST24 all of which take into account the polarized synchrotron data.
@@
In contrast, in the outer Galaxy at high latitudes ($|b|\gtrsim 20^{\circ}$), all coherent field models predict relatively small deflections, likely not exceeding $30^{\circ}$. Thus, if the random deflections are similarly small (see Section II.3), this region may be a promising place to search for UHECR sources. Interestingly, the TA hotspot, marked by the black circle in Fig. 1, is located in this region.
@@ II.2 Deflection directions
In Fig. 5, we present the Amaterasu source localization. The purple contour for the KST24 model was obtained by varying the model parameters, as well as particle energy, within their one-sigma uncertainty. Additionally, $5^{\circ}$ smearing by a toy model of the turbulent GMF was taken into account, following the prescription from [47]. For the KST24 contour, the Amaterasu particle was assumed to be an iron nucleus. The contour size is primarily driven by the observed energy uncertainty. This, however, does not include “systematic” uncertainties related to modeling of GMF, so the actual uncertainties are larger. This is illustrated by the green contour in Fig. 5, which shows the source localization from a Monte Carlo analysis of [43] for the UF23 models. Note that, for both the green contour from [43] and the purple contour presented in this paper for KST24, the particles’ energy was rescaled to $212$ EeV to account for systematic bias in energy reconstruction for heavy primaries. Note also that the observed direction of the Amaterasu event points close to the Loop I, the bright feature in the polarized synchrotron data (shown with black dashed lines in Fig. 5). Recently, it was shown to be a Galactic-scale outflow [48, 49], but it is not included in any GMF model. Incorporating this feature into GMF models could influence the reconstructed direction.
II.3 Deflections in turbulent GMF
In the previous Sections, we discussed UHECR deflections in four different models of the coherent GMF. At the same time, the GMF also has a turbulent component, whose effect on UHECR deflections has been studied extensively in the literature (see, for instance [50, 51].)
A general conclusion from these studies is that the coherent and turbulent components of the GMF affect UHECRs in qualitatively different ways: the coherent GMF shifts the entire source image, while the turbulent GMF produces a halo around the central position determined by the coherent field. Based on the analysis of the RM data, [52, 47] concluded that the typical deflection angle of particle with the rigidity of $R=10$ EV due to the turbulent GMF is of the order of 10 above the Galactic plane. A similar conclusion was reached by [53, 54] in their analysis of the total synchrotron intensity measured by WMAP. If so, the deflection due to turbulent GMF dominate only in a small fraction of the sky, in particular at high Galactic latitudes in the direction of the outer Galaxy, see Fig. 1.
[…]
@@ III Discussion and Summary
In summary, in most directions on the sky, UHECR deflections are still too uncertain — both in direction and magnitude — to enable source identification through backtracking. A notable exception is the region at high latitudes, roughly at $30^{\circ}<b<70^{\circ}$ in the outer Galaxy, where all existing models consistently predict relatively small deflections by the coherent GMF. Interestingly, this is the same region where the TA hotspot has been observed. Regarding the Amaterasu particle, the precise localization of its source cannot be achieved with the KST24 model, although the reconstructed source region is close to that found with the UF23 model.